OpenStudy (kainui):

Show geometrically this is true:

OpenStudy (kainui):

\[\Large \arctan(1)+\arctan(2)+\arctan(3)=\pi\]

OpenStudy (dan815):



OpenStudy (michele_laino):

Please note that if I make the tangent of both sides, I get: \[\begin{gathered} \tan \left[ {\left( {\arctan (1) + \arctan (2)} \right) + \arctan (3)} \right] = \tan \left( \pi \right) \hfill \\ \frac{{\tan \left( {\arctan (1) + \arctan (2)} \right) + 3}}{{1 - \tan \left( {\left( {\arctan (1) + \arctan (2)} \right)} \right) \cdot 3}} = \tan \left( \pi \right) \hfill \\ \frac{{\frac{{1 + 2}}{{1 - 1 \cdot 2}} + 3}}{{1 - \frac{{1 + 2}}{{1 - 1 \cdot 2}} \cdot 3}} = \tan \left( \pi \right) \hfill \\ 0 = 0 \hfill \\ \end{gathered} \]

OpenStudy (dan815):

its gotta be geometrically brooo

OpenStudy (dan815):